OpenStudy (anonymous):

1/x-1+2/x=0?? help solve

OpenStudy (anonymous):

use proper brakets

OpenStudy (anonymous):

\[\frac{ 1 }{ x-1} +\frac{ 2 }{ x } = 0 \] \[\frac{ 1 }{ x-1 } = -\frac{ 2 }{ x }\]

OpenStudy (anonymous):

now cross multiply

OpenStudy (anonymous):

Thanx!

OpenStudy (anonymous):

welcome